OpenStudy (anonymous):

You have taken up gardening for relaxation and have decided to fence in your new rectangular shaped masterpiece. The length of the garden is 2 meters and 28 meters of fencing is required to completely enclose it. What is the width of the garden? a. 24 m b. 12 m c. 56 m d. 14 m

OpenStudy (anonymous):

12 m. 28-2-2=24/2=12